From GlueXWiki
Jump to: navigation, search

January 28, 2016 Drift Chamber meeting


  1. Instructions for Bluejeans meeting connection
  2. Meeting ID: 290664653
  3. To join via a Web Browser, go to the page [1] https://bluejeans.com/290664653.


  1. Collaboration meeting Feb 2016
  2. NIM Papers
  3. Run Readiness Run Coordination Meetings: Spring 2016 Run
  4. Electronics (Fernando, Chris, Nick)
  5. Other

Naomi chaired a separate meeting (with the same people) on the fADC125 firmware: FA125_firmware_check


Participants: Curtis, Naomi (CMU) Eugene, Cody, Luke, Chris, Simon, Beni, and Lubomir (JLab).

Collaboration meeting

- Mike will report on the CDC status.

- Lubomir will give Simon 4-5 slides with the FDC update mostly related to the cathode hit reconstruction, as a part of Simon's track reconstruction talk. There will be no separate FDC talk.

- Segey will report on the TRD results from the last run.

- Naomi will have a talk on the fADC125 in another session but it might be moved to the tracking session (Curtis).

NIM papers

- Mike set up a new ShareLaTex tool for collaborative editing that will be used for the CDC paper (see link above).

- Lubomir is working on a contribution to the VCI conference (middle of February). The contribution will be published as a NIM paper but limited to 4 pages only; the deadline is middle of March. This contribution will be a separate paper from the general FDC paper, and will concentrate on the special studies mostly with the spare package.

Run Readiness

- Beni: the CDC is stable, but he hasn't checked yet the occupancies in all the channels and if some cables are swapped (as it happened before).

- Luke checked the occupancies in all the FDC channels and showed a spreadsheet (linked above) with the results. Wires: in total three channels with no or low signals: all are first channels of the pre-amp cards (in the first package) indicating maybe connection problem. The four wire cards that were noisy before are now OK, but there's a new one with noise. Strips: total number of bad channels about the same - four channels that had problems before are OK, but five new problems all in one cathode in the fourth package.

- Lubomir attempted to fix a HV channel in the most downstream cell 6 of the fourth FDC package. A 10cm Q-tip was needed to reach and clean the HV wire soldering place. First plot above shows the voltage (green, scale not seen, 2000V and 2200V on the plot) and the current (blue, scale in microA). After the cleaning (in the morning of 01/25) the dark current at 2200V went down from unstable ~8 microA to stable 5 microA and then lower. Third plot shows the approximate linearity of the current with the voltage going from 2000V down to 0V. The other plots are zoomed versions of the first plot at different times.